# Unsupervised Anomaly Detection through Mass Repulsing Optimal Transport

### v1

@@ Abstract
Detecting anomalies in datasets is a longstanding problem in machine learning. In this context, anomalies are defined as a sample that significantly deviates from the remaining data. Meanwhile, optimal transport (OT) is a field of mathematics concerned with the transportation, between two probability measures, at least effort. In classical OT, the optimal transportation strategy of a measure to itself is the identity. In this paper, we tackle anomaly detection by forcing samples to displace its mass, while keeping the least effort objective. We call this new transportation problem Mass Repulsing Optimal Transport (MROT). Naturally, samples lying in low density regions of space will be forced to displace mass very far, incurring a higher transportation cost. We use these concepts to design a new anomaly score. Through a series of experiments in existing benchmarks, and fault detection problems, we show that our algorithm improves over existing methods.
@@ 1 Introduction
An anomaly, or an outlier, is a data point that is significantly different from the remaining data (Aggarwal, 2017), to such an extent that it was likely generated by a different mechanism (Hawkins, 1980). From the perspective of machine learning, Anomaly Detection (AD) wants to determine, from a set of examples, which ones are likely anomalies, typically through a score. This problem finds applications in many different fields, such as medicine Salem et al. (2013), cyber-security Siddiqui et al. (2019), and system monitoring Isermann (2006), to name a few. As reviewed in Han et al. (2022), existing techniques for AD are usually divided into unsupervised, semi-supervised and supervised approaches, with an increasing need for labeled data. In this paper, we focus on unsupervised AD, which does not need further labeling effort in constituting datasets. As discussed in Livernoche et al. (2024), the growing number of applications involving high-dimensional and complex data begs the need for non-parametric algorithms.
@@
Meanwhile, Optimal Transport (OT) is a field of mathematics concerned with the transportation of masses at least effort Villani et al. (2009). In its modern treatment, one can conceive transportation problems between probability distributions, which has made an important impact in machine learning research (Montesuma et al., 2024a). Hence, OT is an appealing tool, as it can be estimated non-parametrically from samples from probability distributions. Likewise, the plethora of computational tools for computing OT (Peyré and Cuturi, 2020; Flamary et al., 2021) further stresses its usability.
@@
In this context, the application of OT for AD is not straightforward, as we are interested in analyzing a single probability distribution, rather than 2. This paper puts forth a new OT problem between a distribution and itself, by restricting where a sample can send its mass to. More specifically, we design an exclusion zone, prohibiting samples from keeping its mass, or sending its mass to a small vicinity. Especially, we assume that anomalies lie in low density regions of space, with only a few samples in their vicinity. By restricting the transport of mass in the vicinity of samples, anomalies are naturally forced to send its mass to the high density region, which is assumed to be far away from the anomaly samples. Hence, anomalies will make a larger transportation effort than normal samples, which can find nearby samples outside the exclusion zone. We show a conceptual illustration of our method in Figure.
@@
While OT has been previously used to compare and aggregate signals in the context of AD (Alaoui-Belghiti et al., 2019, 2020), to the best of our knowledge ours is the first general purpose OT-based algorithm for AD. Furthermore, we propose a new OT problem based on the engineering of the ground-cost, which has links to OT with repulsive costs (Di Marino et al., 2017). We benchmark our algorithm in a comprehensive list of datasets, including tabular, computer vision, and natural language processing proposed by Han et al. (2022), besides fault detection (Reinartz et al., 2021; Montesuma et al., 2024b).
@@ 2 Related Work
Anomaly Detection. Following Han et al. (2022), AD methods can be mainly divided into 3 categories. First, supervised methods consider AD through the lens of binary classification under class imbalance. Second, semi-supervised methods either consider partially labeled data, or labeled normal samples, so that an algorithm can be characterize what a normal sample is. The third, more challenging category is unsupervised AD, where the training data contains both anomalies and normal samples and labels are not available. This paper considers precisely the last setting. Next, we review ideas in unsupervised AD.
@@
A fourth kind of emerging ideas in the field considers generative modeling for detecting anomalies. In this context, one devises a model for the data density $P$, for instance, through neural nets Goodfellow et al. (2014); Kingma and Welling (2013). The idea from these models, is that, either the data density can be accurately estimated (Dias et al., 2020), or some score can be derived from the generative model (Livernoche et al., 2023).
@@
In this paper, we consider a different kind of repulsive cost, which we call the mass repulsive cost (see, e.g., Section 3.2 below). Our notion of cost defines an exclusion zone where sending mass is too costly. This exclusion zone is defined on the nearest neighbor analysis of points in a probability measure. As a result, our approach captures the local characteristics of the probability measure being analyzed, especially its density. A special characteristic of our approach is to give a sense of the transportation from a measure to itself.
Optimal Transport-based Anomaly Detection. Previous works (Alaoui-Belghiti et al., 2019, 2020) have considered OT for AD. These works proposed a distance-based detection mechanism, in which isolated samples are considered anomalies. OT contributes to this setting, by defining a rich metric between samples. Especially, these works considered AD in time series data, and OT is used to compute distances between those in the frequency domain, under a Chebyshev ground cost. In comparison with these methods, ours is notably general purpose, that is, we do not assume data to be time series. Instead of using a Chebyshev ground cost, we model the AD problem with a repulsive cost.
@@ 3.1 Optimal Transport
$\hat{P}(\mathbf{x})=\sum_{i=1}^{n}p_{i}\delta(\mathbf{x}-\mathbf{x}_{i}^{(P)}).$ (2)
Since the $p_{i}^{\prime}s$ represent the probability of sampling $\mathbf{x}_{i}^{(P)}$, $\sum_{i}p_{i}=1$ and $p_{i}\geq 0$. Plugging back equation 2 into equation 1 leads to a finite linear program,
$\hat{\gamma}=\underset{\gamma\in\Gamma(\mathbf{p},\mathbf{q})}{\text{arg min}}\,\langle\gamma,\mathbf{C}\rangle_{F}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}\underbrace{c(\mathbf{x}_{i}^{(P)},\mathbf{x}_{j}^{(Q)})}_{C_{ij}},$ (3)
where the optimization variables are the coefficients $\gamma_{ij}$ of the transport plan, and $\gamma\in\Gamma(\mathbf{p},\mathbf{q})$ indicates that $\sum_{i}\gamma_{ij}=q_{j}$ and $\sum_{j}\gamma_{ij}=p_{i}$. As a linear program, the solution $\gamma$ is a sparse matrix with at most $n+m-1$ non-zero elements (Peyré and Cuturi, 2020). Solving it through the celebrated Simplex algorithm (Dantzig, 1983), which has computational complexity $\mathcal{O}(n^{3}\log n)$ and storage complexity $\mathcal{O}(n^{2})$ (i.e., storing each $\gamma_{ij}$).
A faster alternative was introduced by Cuturi (2013), who shown that adding an entropic regularization to equation 3 leads to a problem that can be solved through Sinkhorn’s algorithm Sinkhorn (1967). From a continuous perspective, this is equivalent to penalizing Kullback-Leibler (KL) divergence between $\gamma$, and the trivial coupling $P\otimes Q=P(x)Q(y)$. This regularization term is related to the entropy of $\gamma$ as discussed in (Peyré and Cuturi, 2020, Chapter 4), hence this problem is called entropic OT.
@@
where $\text{KL}(\gamma|\xi)=\int_{\mathcal{X}\times\mathcal{X}}\log\biggr{(}\dfrac{d\gamma}{d\xi}(x,y)\biggr{)}d\gamma(x,y)+\int_{\mathcal{X}\times\mathcal{X}}(d\xi(x,y)-d\gamma(x,y))$ is the KL divergence between measures $\gamma$ and $\xi$.
@@
where $H(\gamma)$ denotes the entropy of the transportation plan. Since equation 5 relies on the Sinkhorn (1967) algorithm rather than linear programming, it has $\mathcal{O}(Ln^{2})$ computational complexity, where $L$ is the number of iterations. In general, the KL and entropic terms in equations 4 and 5 has a smoothing effect over the joint measure $\gamma_{\epsilon}$. As a result, $\hat{\gamma}_{\epsilon}$ has more non-zero elements than $\hat{\gamma}$.
In the next section, we explore a new transportation with a single probability measure. This problem is understood as the transportation of $P$ to itself, when the samples form $P$ are forced to send their mass outside their immediate neighborhood.
@@ 3.2 Mass Repulsing Optimal Transport
In this section, we propose a new OT problem called MROT. This problem is inspired by the Kantorovich formulation, described in section 3.1. However, instead of considering two different probability measures $P,Q$, it considers the transportation of $P$ to itself. Due the properties of OT, if we consider the OT plan $\gamma^{\star}={\text{OT}}(P,P)$, it is supported on the set $\{(x,x),x\in\mathcal{X}\}$, i.e., each point keeps its own mass (Santambrogio, 2015). This motivates our new problem, in which we force points to repell its mass. We do so, by engineering the ground-cost $c:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}_{+}$. For pairs $x\in\mathcal{X}$, $y\in\mathcal{X}$,
$\tilde{c}_{k}(x,y)=\begin{cases}c(x,y)&\text{ if }y\not\in\mathcal{N}_{k}(x),\\ \underset{z\in\mathcal{N}_{k}(x)}{\text{sup}}\,c(x,z)&\text{ otherwise}.\end{cases}$ (6)
@@
this principle is different from repulsive costs (Di Marino et al., 2017), which are designed to model the interaction between particles in multi-marginal OT. Indeed, in our work we are designing a transportation problem from a probability measure to itself. While repulsive costs incentive transportation towards distant points in space, our mass repulsing cost induces transportation just outside an exclusion zone.
While the mass repulsing principle can be applied to any ground-cost $c$, let us focus on the metric case, i.e., when $c(x,y)=d(x,y)^{2}$, where $(\mathcal{X},d)$ is a metric space. The ground-cost we are proposing is essentially defining an exclusion zone, i.e., the $k-$vicinity $\mathcal{N}_{k}(x)$ centered at $x$, where $x$ cannot send its mass. Of course, since $x\in\mathcal{N}_{k}(x)$, it is forced to send its mass elsewhere – thus, we call this idea mass repulsive OT.
@@
Our main hypothesis for anomaly detection is that anomalous points lie in low density regions of $P$. On the one hand, If these points are forced to send its mass outside its vicinity, it will be forced to send it to high density regions of $P$ – otherwise, mass conservation in OT would not hold. On the other hand, if the exclusion zone is smaller than high density regions of $P$, points on these regions will be sent close-by. As a result, anomalous points will have a higher transportation cost, given by $\gamma(x,y)\tilde{c}_{k}(x,y)$, than normal points. We thus consider the following OT problem,
@@
$\gamma^{\star}_{k,\epsilon}$ $=\text{MROT}_{k,\epsilon}(\hat{P})=\underset{\gamma\in\Gamma(\mathbf{p},\mathbf{p})}{\text{min}}\,\langle\gamma,\tilde{\mathbf{C}_{k}}\rangle_{F}-\epsilon H(\gamma).$ (8)
@@
To ground-up intuition about our proposed problem, we show in Figure 1 a comparison between the commonly used squared Euclidean cost, our engineered cost, and a repulsive cost associated with the Coulomb interaction (Di Marino et al., 2017). In traditional OT, sending mass to far away regions of space is costly. As a result, points are encouraged to keep their own mass as close as possible. Consequently, OT from $P$ to itself is the trivial plan $\gamma=\text{Id}$.
Figure 1: Comparison between different ground costs for $x_{0}=(0,0)$. From left to right: Squared euclidean cost $c(x,x_{0})=\lVert x-x_{0}\rVert_{2}^{2}$, Engineered cost (c.f., equation 6) and repulsive cost.
@@
Both our engineered cost $\tilde{c}_{k}$ and the repulsive cost $(1-c(x,x_{0}))^{-1}$ assign high costs to points near $x_{0}$. In particular, within the nearest-neighbor radius $\mathcal{N}{k}(x_{0})$, our engineered cost $\tilde{c}{k}(x,x_{0})$ reaches its maximum. However, there is an important difference between these costs. Our engineered cost promotes transportation to the immediate region outside $\mathcal{N}_{k}$, whereas the repulsive cost drives samples as far as possible from $x_{0}$. As we discussed previously, this feature of the ground-cost allows us to build an anomaly score.

### v2

@@ Abstract
Detecting anomalies in datasets is a longstanding problem in machine learning. In this context, anomalies are defined as a sample that significantly deviates from the remaining data. Meanwhile, Optimal Transport (OT) is a field of mathematics concerned with the transportation, between two probability distribution, at least effort. In classical OT, the optimal transportation strategy of a distribution to itself is the identity, i.e., each sample keeps its mass. In this paper, we tackle anomaly detection by forcing samples to displace its mass, while keeping the least effort objective. We call this new transportation problem Mass Repulsing Optimal Transport (MROT). Naturally, samples lying in low density regions of space will be forced to displace mass very far, incurring a higher transportation cost. In contrast, samples on high density regions are able to send their mass just outside an exclusion zone. We use these concepts to design a new anomaly score. Through a series of experiments in existing benchmarks, and fault detection problems, we show that our algorithm improves over existing methods.
[…]
@@ 1 Introduction
An anomaly, or an outlier, is a data point that is significantly different from the remaining data (Aggarwal, 2017), to such an extent that it was likely generated by a different mechanism (Hawkins, 1980). From the perspective of machine learning, Anomaly Detection (AD) wants to determine, from a set of examples, which ones are likely anomalies, typically through an interpretable score. This problem finds applications in many different fields, such as medicine (Salem et al., 2013), cyber-security (Siddiqui et al., 2019), and system monitoring (Isermann, 2006), to name a few. As reviewed in Han et al. (2022), existing techniques for AD are usually divided into unsupervised, semi-supervised and supervised approaches, with an increasing need for labeled data. In this paper, we focus on unsupervised AD, which does not need further labeling effort in constituting datasets.
@@
Meanwhile, Optimal Transport (OT) is a field of mathematics concerned with the transportation of masses at least effort Villani et al. (2009). In its modern treatment, one can conceptualize transportation problems between probability distributions, which has made an important impact in machine learning research (Montesuma et al., 2024a). Hence, OT is an appealing tool, as it can be estimated non-parametrically from samples from probability distributions. Likewise, the plethora of computational tools for computing OT (Peyré and Cuturi, 2020; Flamary et al., 2021) further stresses its usability.
@@
In this context, the application of OT for AD is not straightforward, as we are interested in analyzing a single probability distribution. We present a new OT problem between a distribution and itself, by restricting where a sample can send its mass to. More specifically, we design an exclusion zone, prohibiting samples from keeping its mass, or sending its mass to a small vicinity. Especially, we assume that anomalies lie in low-density regions of space, with only a few samples in their vicinity. By restricting the transport of mass in the vicinity of samples, anomalies are naturally forced to send their mass to the high-density region, which is assumed to be far away from the anomaly samples. Hence, anomalies will have an overall higher transportation effort than normal samples, which can find nearby samples outside the exclusion zone. We show a conceptual illustration of our method in Figure 1.
@@
Although OT has been previously used to compare and aggregate signals in the context of AD (Alaoui-Belghiti et al., 2019, 2020), to the best of our knowledge ours is the first general purpose OT-based algorithm for AD. Furthermore, we propose a new OT problem based on the engineering of the ground-cost, which has links to OT with repulsive costs (Di Marino et al., 2017). We benchmark our algorithm in a comprehensive list of datasets, including tabular, computer vision, and natural language processing proposed by Han et al. (2022), besides fault detection (Reinartz et al., 2021; Montesuma et al., 2024b).
Figure 1: Mass Repulsive Optimal Transport. Our method engineers the ground-cost $c(x_{i}^{(P)},x_{j}^{(P)})$ between samples of a distribution $P$. That way, samples are forced to send their mass outside a vicinity defined through its nearest neighbors (shaded gray areas). This design leads to a transportation plan, such that anomalous samples incur in an higher transportation effort than normal samples. We use these efforts to build an interpretable and generalizable anomaly score.
@@ 2 Related Work
Anomaly Detection. Following Han et al. (2022), AD methods can be mainly divided into 3 categories. First, supervised methods consider AD through the lens of binary classification under class imbalance. Second, semi-supervised methods either consider partially labeled data, or labeled normal samples, so that an algorithm can characterize what a normal sample is. The third, more challenging category is unsupervised AD, where the training data contains both anomalies and normal samples and labels are not available. This paper considers precisely the last setting. Next, we review ideas in unsupervised AD.
@@
Deep Learning-based Anomaly Detection. As Pang et al. (2021) reviews, deep learning is used for AD in mainly 3 ways: for i) learning powerful feature extractors, ii) learning representations of normality, and iii) building an end-to-end anomaly score. Besides iii, i and ii allow for the synergy between the deep and traditional AD algorithms. In fact, one can use classical AD strategies over the latent space of a neural net. We explore this possibility in Section 4.1.
In addition to this possibility, we give a few examples of how deep learning has contributed to AD. For example, Ruff et al. (2018) proposed a strategy for performing deep one class classification, which echoes the ideas of One Class SVM (OCSVM) Schölkopf et al. (1999).
Furthermore, generative modeling has deeply contributed to AD. For instance, generative adversarial nets (Goodfellow et al., 2014) can serve as a modeling step for capturing features of the underlying distribution $P$. For instance, in Schlegl et al. (2019) and Zenati et al. (2018), the authors combine reconstruction errors with the distance of the predictions of the discriminator for building an anomaly score. Akcay et al. (2018) takes this idea further. This strategy creates a encoder-decoder-encoder architecture for the generator network. The model is trained to model normal images, by mapping them to the latent space, reconstructing back the image, then mapping again to the latent space. Anomalous images are then detected via the usual reconstruction loss, plus the distance between latent representations and the difference between discriminator’s predictions. Another direction consists of using variational inference (Dias et al., 2020) or diffusion models (Livernoche et al., 2024).
@@
In this paper, we consider a different kind of repulsive cost, which we call the mass repulsive cost (see, e.g., Section 3.2 below). Our notion of cost defines an exclusion zone, based on its nearest neighbors, where sending mass is too costly. As a result, our approach captures the local characteristics of the probability distribution being analyzed, especially its density. A special characteristic of our approach is to give a sense of the transportation from a distribution to itself.
Optimal Transport-based Anomaly Detection. Previous works (Alaoui-Belghiti et al., 2019, 2020) have considered OT for AD. These works proposed a distance-based detection mechanism, in which isolated samples are considered anomalies. OT contributes to this setting, by defining a rich metric between samples. Especially, these works considered AD in time series data, and OT is used to compute distances between those time series in the frequency domain, under a Chebyshev ground cost. In comparison with these methods, ours is notably general purpose, that is, we do not assume data to be time series. Instead of using a Chebyshev ground cost, we model the AD problem with a repulsive cost.
@@ 3.1 Optimal Transport
$\hat{P}(\mathbf{x})=\sum_{i=1}^{n}p_{i}\delta(\mathbf{x}-\mathbf{x}_{i}^{(P)}),\,\sum_{i=1}^{n}p_{i}=1,p_{i}\,\geq 1,\forall i.$ (2)
Plugging back equation 2 into equation 1 leads to a finite linear program,
$\hat{\gamma}=\underset{\gamma\in\Gamma(\mathbf{p},\mathbf{q})}{\text{arg min}}\,\langle\gamma,\mathbf{C}\rangle_{F}=\sum_{i=1}^{n}\sum_{j=1}^{m}\gamma_{ij}\underbrace{c(\mathbf{x}_{i}^{(P)},\mathbf{x}_{j}^{(Q)})}_{C_{ij}},\,\Gamma(\mathbf{p},\mathbf{q})=\{\gamma:\sum_{i}\gamma_{ij}=q_{j},\,\sum_{i}\gamma_{ij}=p_{i}\text{ and }\gamma_{ij}\geq 0\}$ (3)
Figure 2: Optimal transport plan $\gamma$ between samples of $P$ (blue) and $Q$ (orange). On the left, we connect samples $(i,j)$ for which $\gamma_{ij}^{\star}>0$. On the right, we show the entries of $\gamma^{\star}$.
where the optimization variables are the coefficients $\gamma_{ij}$ of the transport plan. Problem 3 is a finite linear program, hence the solution $\gamma$ is a sparse matrix with at most $n+m-1$ non-zero elements (Peyré and Cuturi, 2020). Solving it through the celebrated Simplex algorithm (Dantzig, 1983), which has computational complexity $\mathcal{O}(n^{3}\log n)$ and storage complexity $\mathcal{O}(n^{2})$ (i.e., storing each $\gamma_{ij}$). We show an example of empirical OT in Figure 2.
A faster alternative was introduced by Cuturi (2013), who shown that adding an entropic regularization to equation 3 leads to a problem that can be solved through Sinkhorn’s algorithm (Sinkhorn, 1967). From a continuous perspective, this is equivalent to penalizing Kullback-Leibler (KL) divergence between $\gamma$, and the trivial coupling $P\otimes Q=P(x)Q(y)$. This regularization term is related to the entropy of $\gamma$ as discussed in (Peyré and Cuturi, 2020, Chapter 4), hence this problem is called entropic OT. Next, we define the entropic OT problem,
@@
where $\text{KL}(\gamma|\xi)=\int_{\mathcal{X}\times\mathcal{X}}\log\biggr{(}\dfrac{d\gamma}{d\xi}(x,y)\biggr{)}d\gamma(x,y)+\int_{\mathcal{X}\times\mathcal{X}}(d\xi(x,y)-d\gamma(x,y))$ is the KL divergence between measures $\gamma$ and $\xi$ (Peyré and Cuturi, 2020, Eq. 4.10).
@@
where $H(\gamma)$ denotes the entropy of the transportation plan. Since equation 5 relies on the Sinkhorn (1967) algorithm rather than linear programming, it has $\mathcal{O}(Ln^{2})$ computational complexity, where $L$ is the number of iterations. In general, the KL and entropic terms in equations 4 and 5 have a smoothing effect on the transportation plan $\gamma_{\epsilon}$. As a result, $\hat{\gamma}_{\epsilon}$ has more non-zero elements than $\hat{\gamma}$.
In the next section, we explore a new transportation problem with a single probability distribution. This problem is understood as the transportation of $P$ to itself, when the samples form $P$ are forced to send their mass outside of their immediate neighborhood.
@@ 3.2 Mass Repulsing Optimal Transport
In this section, we propose a new OT problem called MROT. This problem is inspired by the Kantorovich formulation, described in section 3.1. However, instead of considering two different probability distributions $P$, $Q$, it considers the transport of $P$ to itself. Due to the properties of OT, if we consider the OT plan $\gamma^{\star}={\text{OT}}(P,P)$, it is supported in the set $\{(x,x),x\in\mathcal{X}\}$, that is, each point keeps its own mass (Santambrogio, 2015). This motivates our new problem, in which we force points to repell its mass. Henceforth, we assume that $c$ comes from a metric $(\mathcal{X},d)$, i.e., $c(x,y)=d(x,y)^{p}$, $p\in[1,+\infty)$. For pairs $x\in\mathcal{X}$, $y\in\mathcal{X}$, and $L\in[0,+\infty)$,
$\tilde{c}(x,y)=\begin{cases}c(x,y)&\text{ if }y\not\in\mathcal{N}(x),\\ L&\text{ otherwise},\end{cases}$ (6)
where $\mathcal{N}(x)$ denotes the vicinity of $x$ (e.g., $k-$nearest neighbors, or an $\rho-$ball centered at $x$). In the next remark, we give some geometric intuition behind the cost engineering, as well as some motivation for the choice of $L$. More generally, in the next section we give a theoretically motivated choice for $L$.
Figure 3: Comparison between different ground costs for $x_{0}=(0,0)$. From left to right: Squared euclidean cost $c(x,x_{0})=\lVert x-x_{0}\rVert_{2}^{2}$, Engineered cost (c.f., equation 6) for two choices of $L$, and repulsive cost.
Remark 3.1.
(Geometric intuition) Assume $x_{0}\in\mathcal{X}=[-2,2]^{2}$ fixed. Our cost engineering strategy replaces the base ground cost $c(x_{0},y)$, in the neighborhood $\mathcal{N}(x_{0})$ with a value $L$. Assume, for argumentation purposes, that $L\rightarrow+\infty$. This choice means that it is infinitely costly for each point to keep its own mass. As a result, samples are encouraged to send their mass to the immediate outside of their neighborhood – thus, we call this idea mass repulsive OT.
To ground some intuition, we show, in Figure 3, a comparison using the squared Euclidean distance as the ground-cost, i.e., $c(x,y)=\lVert x-y\rVert_{2}^{2}$. Alongside the base cost, we show our engineered cost, and the repulsive cost $(1+c(x,y))^{-1}$. Here, we define the neighborhood of $x_{0}$ through the $\rho-$closed ball centered at $x_{0}$, i.e., $\mathcal{N}(x_{0})=\{y:\mathbb{R}^{2}:\lVert x-y\rVert_{2}\leq\rho\}$. Likewise, we use $\rho=1$, and, $L=\text{sup}_{z\in\mathcal{N}(x_{0})}c(x,z)$. This particular choice leads to nice theoretical properties for the continuous MROT problem (see equation 7 below).
Due to the particular choices we made, the supremum takes its value on the frontier $\partial\mathcal{N}(x_{0})=\{z:\mathbb{R}^{2}:\lVert x_{0}-z\rVert_{2}=\rho\}$, in which case $c(x_{0},z)=\rho^{2}$, independently of $x_{0}$. This choice for $L$ effectively flattens the ground cost around the vicinity of $x_{0}$. As we show in the next section, this is essential for ensuring the existence of the continuous MROT problem. In practice, it is beneficial to set $L$ to a large, finite constant to ensure points are transported outside $\mathcal{N}(x_{0})$. Note that, as we remark in the next section, this does not harm the existence of the discrete MROT solution. We show an example of this choice by setting $L=\text{sup}_{z\in\mathcal{X}}\lVert x-z\rVert_{2}^{2}$.
@@
Our principle of mass repulsion is different from repulsive costs (Di Marino et al., 2017), which are designed to model the interaction between particles in multimarginal OT (Pass, 2015). Indeed, we design a transportation problem from a probability distribution to itself. Hence, while repulsive costs incentive transportation towards distant points in space, our mass repulsing cost induces transportation just outside an exclusion zone. Henceforth, we focus on $p=2$ for the Euclidean distance. The ground-cost we propose essentially defines an exclusion zone around points $\mathbf{x}$, where these points are discouraged from sending their mass.
@@
Our main hypothesis for anomaly detection is that anomalous points lie in low density regions of $P$. On the one hand, If these points are forced to send its mass outside its vicinity, it will be forced to send it to high density regions of $P$ – otherwise, mass conservation in OT would not hold. On the other hand, if the exclusion zone is smaller than high density regions of $P$, points on these regions will be sent close-by. As a result, anomalous points will have a higher transportation cost (c.f., equations 9 and 10 below) than normal points. We thus consider the following OT problem,
@@
$\gamma^{\star}_{\epsilon}$ $=\text{MROT}_{\epsilon}(\hat{P})=\underset{\gamma\in\Gamma(\mathbf{p},\mathbf{p})}{\text{min}}\,\langle\gamma,\tilde{\mathbf{C}}\rangle_{F}-\epsilon H(\gamma).$ (8)
@@
Considering again Figure 3, we focus on comparing the commonly used squared Euclidean cost, our engineered cost, and a repulsive cost associated with the Coulomb interaction (Di Marino et al., 2017). In traditional OT, sending mass to distant regions of space is costly. As a result, points are encouraged to keep their own mass as close as possible. Consequently, OT from $P$ to itself is the trivial plan $\gamma=\text{Id}$.
@@
For our engineered cost, we use $\mathcal{N}(\mathbf{x})=\{\mathbf{y}:\lVert\mathbf{x}-\mathbf{y}\rVert_{2}^{2}\leq\rho\}$, $\rho=1$. Both our engineered cost $\tilde{c}$ and the repulsive cost $(1-c(x,x_{0}))^{-1}$ assign high costs to points near $x_{0}$. In particular, within the neighborhood $\mathcal{N}(x_{0})$, our engineered cost $\tilde{c}(x,x_{0})$ reaches its maximum. Hence, there is an important difference between MROT and OT with repulsive costs. Our engineered cost promotes transportation to the immediate region outside $\mathcal{N}$, whereas the repulsive cost drives samples as far as possible from $x_{0}$. As we discussed previously, this feature of the ground-cost allows us to build an anomaly score.
3.3 Theoretical Analysis
In this section, we provide a proof for the existence of MROT plans. While at least one solution of equation 7 exists, its uniqueness is not, in general, guaranteed. In this section, we denote the set of measures over a set $\mathcal{X}$ by $\mathcal{P}(\mathcal{X})$, and the power-set of $\mathcal{X}$ by $\mathbb{P}(\mathcal{X})$.
From classical OT theory, the Kantorovich problem (c.f., equation 1) has a solution under mild conditions on the ambient space $\mathcal{X}$ and the ground-cost $c$, see Santambrogio (2015). Indeed, $\mathcal{X}$ is required to be Polish (i.e., complete, separable metric space) and $c$ must be lower semi-continuous. For simplicity, we state our results on $d-$dimensional Euclidean spaces, which are the setting for our experiments. We start by re-stating (Santambrogio, 2015, Theorem 1.5),
Theorem 3.1.
(Existence of Optimal Transport plans) Let $\mathcal{X}$ be a Polish space, $P$, and $Q\in\mathcal{P}(\mathcal{X})$, and $c:\mathcal{X}\times\mathcal{X}\rightarrow[0,+\infty]$ be lower semi-continuous. Then equation 1 admits a solution.
Since we are dealing with $d-$dimensional Euclidean spaces, the condition on the ambient space is already satisfied. It remains the question of the lower semi-continuity of our engineered cost. In the next theorem, we prove that, for the squared Euclidean cost, a solution for MROT exists,
Theorem 3.2.
Let $c(x,y)=\lVert x-y\rVert_{2}^{2}$ and $\mathcal{N}(x)=\{y\in\mathbb{R}^{d}:\lVert x-y\rVert_{2}\leq\rho\}$. Then the continuous MROT problem (equation 7) admits a solution.
Proof.
Let $x_{0}\in\mathbb{R}^{d}$ and $\rho>0$ be given. Our proof relies on the decomposition of $y\in\mathbb{R}^{d}$ in 3 regions: i) $\lVert x_{0}-y\rVert_{2}>\rho$, ii) $\lVert x_{0}-y\rVert_{2}<\rho$, and iii) $\lVert x-y\rVert_{2}=\rho$. For i) and ii), $\tilde{c}$ is a continuous function of its inputs. Indeed, for $\lVert x_{0}-y\rVert_{2}>\rho$, $\tilde{c}(x,y)=\lVert x_{0}-y\rVert_{2}^{2}$. Likewise, for $\lVert x_{0}-y\rVert_{2}<\rho$, we have $\tilde{c}(x,y)=\rho^{2}$. Now, we need to prove that $\tilde{c}$ is continuous on $\lVert x_{0}-y\rVert_{2}=\rho$. We sub-divide this into 2 cases. First, assume we have a sequence $(y_{n})$, $y_{n}\rightarrow y$, $y_{n}\not\in\mathcal{N}(x)$. Since $y_{n}\not\in\mathcal{N}(x_{0})$, we have $\tilde{c}(x_{0},y_{n})=\lVert x_{0}-y_{n}\rVert_{2}^{2}$. From the continuity of the squared Euclidean distance, $\tilde{c}(x_{0},y_{n})\rightarrow\lVert x_{0}-y\rVert_{2}^{2}=\rho^{2}$. Second, assume we have a sequence $(y_{n}^{\prime}),y_{n}^{\prime}\rightarrow y$, $y_{n}^{\prime}\in\mathcal{N}(x_{0})$. Then, $\tilde{c}(x_{0},y_{n}^{\prime})=\rho^{2}\rightarrow\tilde{c}(x_{0},y)=\rho^{2}$. Hence, $\tilde{c}$ is continuous. ∎
The principles used in the previous theorem can be generalized to other costs and neighborhoods, as long as: i) the base cost $c(x,y)$ is continuous, ii) the neighborhood $\mathcal{N}(x)$ is compact, and iii) the supremum $\text{sup}_{z\in\mathcal{N}(x)}c(x,z)$ is attained in the boundary of $\mathcal{N}(x)$, $\forall x\in\mathcal{X}$. The bottom line of our analysis is that, in general, one needs to carefully engineer the ground cost, through the choice of the base ground cost $c$, and the neighborhood function $\mathcal{N}$.
Although existence may hold, the fact that the engineered cost introduces a flat region on the neighborhood $\mathcal{N}(x)$ of $x$ makes it difficult to establish the uniqueness of the MROT plan. We recall from (Figalli and Glaudo, 2021, Section 2.6) that, in classical OT, the existence of a unique transport plan is related to the existence of a Monge map between $P$ and $Q$. From (Figalli and Glaudo, 2021, 2.7.1), this means that we need 3 conditions: i) $x\mapsto\tilde{c}(x,y)$ is differentiable, ii) $y\mapsto\nabla_{x}\tilde{c}(x,y)$ is injective, and iii) for $\rho>0$, and $B_{\rho}=\{x\in\mathcal{X}:\lVert x\rVert\leq\rho\}$, $|\nabla\tilde{c}(x,y)|\leq C_{\rho}$ for every $x\in B_{\rho}$. As it turns out, neither of these conditions holds for our engineered cost, so the MROT plan is likely not unique.
Now, let us focus on the discrete case, which is of practical interest to us. Note that $\Gamma(\mathbf{p},\mathbf{p})$ is non-empty as $\gamma_{ij}=p_{i}p_{j}\in\Gamma(\mathbf{p},\mathbf{p})$. Furthermore, this set is compact, and the objective in equation 8 is continuous. Then, by Weierstrass theorem (Santambrogio, 2015, Box 1.1), there exists a minimizer to the discrete setting. Here, note that we did not have to impose continuity or lower semi-continuity on the neighborhood function $\mathcal{N}$ or the ground cost $\tilde{c}$. As in classical OT, the solution to this problem is not unique, since the objective function is not strictly convex. Following this line, adding the entropic penalty (i.e., $\epsilon>0$ in equation 8) makes the objective strictly convex (Peyré and Cuturi, 2020, Section 4.1), which guarantees the uniqueness of the OT plan, and hence, the uniqueness of the MROT plan.